OpenStudy (anonymous):

I need help with the steps with this problem as well, log(z-3)=2

jimthompson5910 (jim_thompson5910):

Assuming you mean log base 10 \[\large \log(z-3)=2\] \[\large z-3=10^2\] \[\large z-3=100\] \[\large z=100+3\] \[\large z=103\]

OpenStudy (anonymous):

In the next problem i'm given a decimal is that the same format. 

jimthompson5910 (jim_thompson5910):

So the problem is \[\large 7\log_{4}(0.6x)=12\] ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\large 7\log_{4}(0.6x)=12\] \[\large \log_{4}(0.6x)=\frac{12}{7}\] \[\large 0.6x=4^{\frac{12}{7}}\] \[\large x=\frac{4^{\frac{12}{7}}}{0.6}\] \[\large x\approx 17.9453359017648\]